solveUse the data in the following table to calculate the payoff and the profits for investments in each of the following July exp
Answer #1

ANSWER - CURRENT PRICE = $150

call payoff = Max[0, spot price - strike price]

put payoff = Max[0, strike price - spot price]

A. CALL OPTION X = $145

call payoff = Max[0, spot price - strike price]

= Max[0, 150- 145]

= 5

loss = cost of premium - payoff

=5.18 - 5

= 0.18 per share loss

B. PUT OPTION AT X = $145

put payoff = Max[0, strike price - spot price]

= max[0, 145 - 150]

= 0

loss = cost of premium - premium

= 0.48 - 0

loss = 0.48 per share

C. CALL OPTION AT X = $150

call payoff = Max[0, spot price - strike price]

= Max[0, 150- 150]

= 0

loss = cost of premium - payoff

= 1.85 - 0

= 1.85 per share loss

D. PUT OPTION AT X = $150

put payoff = Max[0, strike price - spot price]

= max[0, 150 - 150]

= 0

loss = cost of premium - premium

=1.81 - 0

loss = 1.81 per share

E. CALL OPTION X = $155

call payoff = Max[0, spot price - strike price]

= Max[0, 150- 155]

= 0

loss = cost of premium - payoff

=0.79 - 0

= 0.79 per share loss

F. PUT OPTION AT X = $155

put payoff = Max[0, strike price - spot price]

= max[0, 155 - 150]

= 5

loss = cost of premium - premium

=5.95 - 5

loss = 0.95 per share